# Skip Tuning: Pre-trained Vision-Language Models are Effective and Efficient Adapters Themselves

### v1

@@ 3.2 Experimental Results
Domain Generalization. The domain generalization setting assesses whether models trained on a source domain can generalize to unseen/target domains. In line with those comparison approaches, we consider the ImageNet dataset as the source domain and the other four ImageNet variants as target domains. Table 2 presents the obtained domain generalization results of Skip Tuning and the state-of-the-art methods. As shown, compared to the previous state-of-the-art method PromptSRC, Skip Tuning achieves time efficiency, and memory efficiency, while establishing 1.5% and 0.55% ACC improvements on the source and target domains, respectively. Besides, Skip Tuning consistently outperforms those competitors on the four target domains without compromising the performance of the tuned model on the source domain, which proves the effectiveness of our Skip Tuning scheme in improving the robustness of the tuned model to domain shifts.
@@ 3.3 Ablation Studies
In this section, we conduct ablative studies to further scrutinize our devised Skip Tuning method. Unless stated otherwise, the average base-to-new generalization performance over the 11 datasets is reported.
@@ 4 Related Work
Prompt Tuning. Prompt tuning (a.k.a. context optimization [42]) has emerged as a parameter-efficient learning paradigm to adapt powerful VLMs to downstream tasks [38, 37, 9, 8]. As a representative method, CoOp [42] enables task adaptation by optimizing a set of prompt vectors within CLIP’s language branch. While effective, CoOp often suffers from limited generalization on new tasks due to overfitting to the base task. A series of schemes are proposed to cope with this problem, e.g. CoCoOp[41], KgCoOp[32], ProGrad[44], TCP [33], and DePT[39]. By incorporating trainable prompts into both the image and text branches of CLIP, MaPLe[17], PromptSRC[18], and CoPrompt[27] demonstrate remarkable performance on both base and new tasks. Despite the great advantages, we reveal that freezing the parameters of VLMs during learning the context vectors neither facilitates the generalization of pre-trained knowledge nor significantly improves memory and time efficiency. Our goal in this work is to adapt VLMs to downstream tasks in an effective and efficient manner.

### v2

@@ 3.2 Experimental Results
Domain Generalization. The domain generalization setting assesses whether models trained on a source domain can generalize to unseen/target domains. In line with those comparison methods, we consider the ImageNet as the source domain and the other four ImageNet variants as target domains. As shown in Table 6, compared to the previous state-of-the-art method PromptSRC, Skip Tuning achieves time efficiency, and memory efficiency, while establishing 1.5% and 0.55% ACC improvements on the source and target domains, respectively. Besides, Skip Tuning consistently outperforms those competitors on the four target domains without compromising the performance of the tuned model on the source domain, which proves the effectiveness of our Skip Tuning scheme in improving the robustness of the tuned model to domain shifts.
@@ 3.3 Ablation Studies
In this section, we conduct ablative studies to further scrutinize our devised Skip Tuning method.
@@ 4 Related Work
Prompt Tuning. Prompt tuning (a.k.a. context optimization [42]) has emerged as a parameter-efficient learning paradigm to adapt powerful VLMs to downstream tasks [38, 37, 9, 7]. As a representative method, CoOp [42] enables task adaptation by optimizing a set of prompt vectors within CLIP’s language branch. While effective, CoOp often suffers from limited generalization on new tasks due to overfitting to the base task. A series of schemes are devised to tackle this problem, e.g. CoCoOp[41], KgCoOp[32], ProGrad[44], TCP [33], and DePT[39]. By adding trainable prompts into both the image and text branches of CLIP, MaPLe[17], PromptSRC[18], and CoPrompt[27] demonstrate remarkable performance on both base and new tasks. Despite the advantages, we reveal that freezing the parameters of VLMs during learning the context vectors neither facilitates the generalization of pre-trained knowledge nor significantly improves memory and time efficiency.
@@ 5 Conclusion
Acknowledgements. This study is supported by grants from the National Natural Science Foundation of China (Grant No. 62425208, No. 62122018, No. 62020106008), the Postdoctoral Fellowship Program of CPSF (Grant No. GZB20240625), the Science and Technology Innovation Committee of Shenzhen Municipality Foundation (Grant No.JCYJ20240813114208012), and Kuaishou Tech.
Supplementary Materials
Appendix A Experimental Setups for FT and PT
While many Prompt Tuning (PT) schemes have reported improved performance and efficiency over the full fine-tuning (FT) baseline, the discrepancy of the implementation details among those schemes obscures the actual performance enhancement. To comprehensively evaluate the progress established by PT, in Section 2.2 of the paper, we perform a comparison between the representative PT method CoOp [42] with the FT baseline in terms of i) the number of learnable parameters, ii) memory usage, iii) time cost, and iv) base-to-new generalization results. The experimental setups of FT and PT are as follows.
The FT baseline involves updating all parameters of the pre-trained CLIP model. For fair comparisons, we maintain the same batch size of 4 and a total of 10 training epochs (as CoOp), the learning rate is set to 1e-5.
Appendix B Cross-Dataset / Domain General. Results
| CoOp | CoCoOp | ProGrad | KgCoOp | MaPLe | ProSRC. | TCP | DePT | CoPro.* | SkipT. | ||
| Source | ImgNet | 71.51 | 71.02 | 72.24 | 70.66 | 70.72 | 71.27 | 71.40 | 72.77 | 72.53 | 72.77 |
| Average | 63.88 | 65.74 | 62.71 | 65.51 | 66.30 | 65.81 | 66.29 | 65.55 | 64.81 | 67.00 | |
| Target | Caltech | 93.70 | 94.43 | 91.52 | 93.92 | 93.53 | 93.60 | 93.97 | 94.23 | 94.37 | 93.43 |
| Pets | 89.14 | 90.14 | 89.64 | 89.83 | 90.49 | 90.25 | 91.25 | 90.03 | 89.50 | 90.10 | |
| Cars | 64.51 | 65.32 | 62.39 | 65.41 | 65.57 | 65.70 | 64.69 | 65.57 | 65.57 | 65.37 | |
| Flowers | 68.71 | 71.88 | 67.87 | 70.01 | 72.23 | 70.25 | 71.21 | 70.57 | 69.63 | 71.97 | |
| Food101 | 85.30 | 86.06 | 85.40 | 86.36 | 86.20 | 86.15 | 86.69 | 86.37 | 85.03 | 86.17 | |
| Aircraft | 18.47 | 22.94 | 20.16 | 22.51 | 24.74 | 23.90 | 23.45 | 23.27 | 23.47 | 25.13 | |
| SUN397 | 64.15 | 67.36 | 62.47 | 66.16 | 67.01 | 67.10 | 67.15 | 66.67 | 67.13 | 67.33 | |
| DTD | 41.92 | 45.73 | 39.42 | 46.35 | 46.49 | 46.87 | 44.35 | 44.97 | 44.33 | 48.00 | |
| EuroSAT | 46.39 | 45.37 | 43.46 | 46.04 | 48.06 | 45.50 | 51.45 | 43.53 | 40.87 | 54.27 | |
| UCF101 | 66.55 | 68.21 | 64.29 | 68.50 | 68.69 | 68.75 | 68.73 | 69.30 | 68.17 | 68.23 | |
| CoOp | CoCoOp | ProGrad | KgCoOp | MaPLe | ProSRC. | DePT | CoPro.* | SkipT. | ||
| Source | ImgNet | 71.51 | 71.02 | 72.24 | 70.66 | 70.72 | 71.27 | 72.77 | 72.53 | 72.77 |
| Average | 59.28 | 59.90 | 59.07 | 60.11 | 60.26 | 60.65 | 58.97 | 60.02 | 61.20 | |
| Target | ImgNet-V2 | 64.20 | 64.07 | 64.73 | 64.10 | 64.07 | 64.35 | 64.70 | 65.40 | 65.67 |
| ImgNet-S | 47.99 | 48.75 | 47.61 | 48.97 | 49.15 | 49.55 | 47.73 | 49.90 | 49.73 | |
| ImgNet-A | 49.71 | 50.63 | 49.39 | 50.69 | 50.90 | 50.90 | 48.33 | 47.53 | 51.13 | |
| ImgNet-R | 75.21 | 76.18 | 74.58 | 76.70 | 76.98 | 77.80 | 75.10 | 77.23 | 78.27 | |
| Time (s) | 31632 | 93917 | 56223 | 31636 | 4942 | 50091 | 22796 | 19161 | 1374 | |
| Cost | Mem. (M) | 15412 | 13622M | 15412 | 15254 | 13786 | 14107 | 13783 | 14131 | 925 |
To evaluate the robustness of different methods to distribution and domain shifts, we conduce comprehensive comparisons in cross-dataset generalization and domain generalization benchmarks.
Cross-Dataset Generalization. The cross-dataset generalization setting assesses whether models trained on a source dataset/distribution can generalize to unseen target datasets/distributions. We follow the common setup of those comparison methods to use ImageNet as the source dataset and the other 10 datasets as target datasets. Table 6 presents the obtained cross-dataset generalization results of our Skip Tuning and other state-of-the-art methods on the source and target datasets. As shown, Skip Tuning achieves the highest overall performance while maintaining the lowest time and memory costs, especially on challenging datasets with significant differences from the source domain (e.g., Aircraft, DTD, EuroSAT). This demonstrates the effectiveness of our Skip Tuning scheme for improving the robustness of the tuned model to distribution shifts.
Domain Generalization. The domain generalization setting assesses whether models trained on a source domain can generalize to unseen/target domains. In line with those comparison approaches, we consider the ImageNet dataset as the source domain and the other four ImageNet variants as target domains. Table 7 reports the obtained domain generalization results of our Skip Tuning and other state-of-the-art methods on the source and target ddomains. As can be seen, our Skip Tuning outperforms other methods in terms of overall performance and maintains the lowest time and memory costs. Besides, Skip Tuning consistently outperforms those competitors on the four target domains without compromising the performance of the tuned model on the source domain, which proves the effectiveness of our Skip Tuning scheme in improving the robustness of the tuned model to domain shifts.
Appendix C Comparisons with Adapter-based Methods
| Method | Cache | LR | BS | Epochs | |
| LoRA | - | 1e-3 | 4 | 10 | - |
| CLIP-adapter | ✗ | 2e-3 | 32 | 100 | - |
| CLIP-adapter† | ✓ | 3.5e-3 | 4 | 10 | - |
| SkipT.(Ours) | ✓ | 2e-5 | 4 | 10 | 6 |
| SkipT.(Ours)† | ✓ | 2e-5 | 4 | 10 | 10 |
| Shot | Dataset | CoOp | CoCoOp | MaPLe | ProSRC. | SkipT. |
| (IJCV’22) | (CVPR’22) | (CVPR’23) | (ICCV’23) | (Ours) | ||
| ImgNet | 66.33 | 69.43 | 62.67 | 68.13 | 69.20 | |
| Caltech | 92.60 | 93.83 | 92.57 | 93.67 | 93.87 | |
| Pets | 90.37 | 91.27 | 89.10 | 92.00 | 91.60 | |
| Cars | 67.43 | 67.22 | 66.60 | 69.40 | 69.63 | |
| Flowers | 77.53 | 72.08 | 83.30 | 85.93 | 83.63 | |
| Food101 | 84.33 | 85.65 | 80.50 | 84.87 | 85.67 | |
| Aircraft | 21.37 | 12.68 | 26.73 | 27.67 | 29.93 | |
| SUN397 | 66.77 | 68.33 | 64.77 | 69.67 | 69.10 | |
| DTD | 50.23 | 48.54 | 52.13 | 56.23 | 54.50 | |
| EuroSAT | 54.93 | 55.33 | 71.80 | 73.13 | 72.23 | |
| UCF101 | 71.23 | 70.30 | 71.83 | 74.80 | 75.30 | |
| Average | 67.56 | 66.79 | 69.27 | 72.32 | 72.24 | |
| Time (s) | 327 | 678 | 911 | 999 | 106 | |
| 1 | Mem. (M) | 4551 | 2888 | 3061 | 3373 | 528 |
| ImgNet | 67.07 | 69.78 | 65.10 | 69.77 | 70.23 | |
| Caltech | 93.07 | 94.82 | 93.97 | 94.53 | 95.10 | |
| Pets | 89.80 | 92.64 | 90.87 | 92.50 | 92.43 | |
| Cars | 70.50 | 68.37 | 71.60 | 73.40 | 73.17 | |
| Flowers | 87.33 | 75.79 | 88.93 | 91.17 | 90.33 | |
| Food101 | 84.40 | 86.22 | 81.47 | 85.70 | 85.90 | |
| Aircraft | 26.20 | 15.06 | 30.90 | 31.70 | 33.53 | |
| SUN397 | 66.53 | 69.03 | 67.10 | 71.60 | 70.60 | |
| DTD | 53.60 | 52.17 | 55.50 | 59.97 | 58.83 | |
| EuroSAT | 65.17 | 46.74 | 78.30 | 79.37 | 78.50 | |
| UCF101 | 73.43 | 73.51 | 74.60 | 78.50 | 78.37 | |
| Average | 70.65 | 67.65 | 72.58 | 75.29 | 75.18 | |
| Time (s) | 561 | 1346 | 1785 | 1947 | 165 | |
| 2 | Mem. (M) | 4551 | 2888 | 3061 | 3373 | 528 |
| ImgNet | 68.73 | 70.39 | 67.70 | 71.07 | 71.40 | |
| Caltech | 94.40 | 94.98 | 94.43 | 95.27 | 95.60 | |
| Pets | 92.57 | 92.81 | 91.90 | 93.43 | 93.33 | |
| Cars | 74.47 | 69.39 | 75.30 | 77.13 | 77.60 | |
| Flowers | 92.17 | 78.40 | 92.67 | 93.87 | 94.27 | |
| Food101 | 84.47 | 86.88 | 81.77 | 86.17 | 86.00 | |
| Aircraft | 30.83 | 24.79 | 34.87 | 37.47 | 39.90 | |
| SUN397 | 69.97 | 70.21 | 70.67 | 74.00 | 73.07 | |
| DTD | 58.70 | 55.04 | 61.00 | 65.53 | 65.70 | |
| EuroSAT | 70.80 | 65.56 | 84.50 | 86.30 | 83.40 | |
| UCF101 | 77.10 | 74.82 | 78.47 | 81.57 | 82.53 | |
| Average | 74.02 | 71.21 | 75.37 | 78.35 | 78.44 | |
| Time (s) | 1036 | 2684 | 3541 | 3841 | 280 | |
| 4 | Mem. (M) | 4551 | 2888 | 3061 | 3373 | 528 |
| ImgNet | 70.63 | 70.63 | 70.30 | 72.33 | 72.40 | |
| Caltech | 94.37 | 95.04 | 95.20 | 95.67 | 95.90 | |
| Pets | 91.27 | 93.45 | 92.57 | 93.50 | 93.40 | |
| Cars | 79.30 | 70.44 | 79.47 | 80.97 | 82.33 | |
| Flowers | 94.97 | 84.30 | 95.80 | 96.27 | 96.60 | |
| Food101 | 82.67 | 86.97 | 83.60 | 86.90 | 86.73 | |
| Aircraft | 39.00 | 26.61 | 42.00 | 43.27 | 46.50 | |
| SUN397 | 71.53 | 70.84 | 73.23 | 75.73 | 74.77 | |
| DTD | 64.77 | 58.89 | 66.50 | 69.87 | 69.77 | |
| EuroSAT | 78.07 | 68.21 | 87.73 | 88.80 | 86.57 | |
| UCF101 | 80.20 | 77.14 | 81.37 | 84.30 | 85.60 | |
| Average | 76.98 | 72.96 | 78.89 | 80.69 | 80.96 | |
| Time (s) | 1988 | 5358 | 7060 | 7636 | 508 | |
| 8 | Mem. (M) | 4551 | 2888 | 3061 | 3373 | 528 |
| ImgNet | 71.87 | 70.83 | 72.33 | 73.17 | 73.83 | |
| Caltech | 95.57 | 95.16 | 96.00 | 96.07 | 96.77 | |
| Pets | 91.87 | 93.34 | 92.83 | 93.67 | 94.00 | |
| Cars | 83.07 | 71.57 | 83.57 | 83.83 | 87.43 | |
| Flowers | 97.07 | 87.84 | 97.00 | 97.60 | 98.17 | |
| Food101 | 84.20 | 87.25 | 85.33 | 87.50 | 87.13 | |
| Aircraft | 43.40 | 31.21 | 48.40 | 50.83 | 55.57 | |
| SUN397 | 74.67 | 72.15 | 75.53 | 77.23 | 76.80 | |
| DTD | 69.87 | 63.04 | 71.33 | 72.73 | 74.07 | |
| EuroSAT | 84.93 | 73.32 | 92.33 | 92.43 | 91.57 | |
| UCF101 | 82.23 | 78.14 | 85.03 | 86.47 | 87.70 | |
| Average | 79.89 | 74.90 | 81.79 | 82.87 | 83.91 | |
| Time (s) | 3879 | 10694 | 14070 | 14403 | 962 | |
| 16 | Mem. (M) | 4551 | 2888 | 3061 | 3373 | 528 |
Apart from Prompt Tuning (PT), adapter-based methods can be used to transfer large pre-trained vision-language models to downstream tasks. Table 8 compares Skip Tuning with representative adapter-based methods. The details of those comparison methods are listed in Table 8.
LoRA: The default implementation of LoRA [15] that adds learnable weights to different blocks of the frozen CLIP model. We adopt a learning rate of 1e-3, a batch size of 4, and training epochs of 10.
CLIP-adapter. The original implementation of CLIP-adapter [10] that attends the learnable adapter module to the frozen CLIP model. The learning rate is 2e-3, the batch size is 32, the number of training epochs is 100.
CLIP-adapter†. A computational efficiency-optimized version of CLIP-adapter [10]. Concretely, the CLIP’s last-layer features are cached once before training and subsequently used to update the weights of the adapter module. We adjust the learning rate to 3.5e-3, the batch size to 4, and the number of training epochs to 10.
Skip Tuning†. A computational efficiency-optimized version of our Skip Tuning, which sets to filter out more shallow features during the LSkip step. This variant is used to investigate the performance of Skip Tuning under extreme efficiency constraints.
Appendix D Few-shot Learning Results
Table 9 presents the few-shot learning performance of different methods across 11 datasets with varying shots of {1, 2, 4, 8, 16}. As can be observed, our Skip Tuning demonstrates superior performance compared to other methods, with significant efficiency advantages in memory usage and time cost across 1-shot to 16-shot settings. Additionally, while PromptSRC and MaPLe enhance CoOp’s performance, both methods incur increased time cost. For instance, PromptSRC[18] improves the average ACC by 2.98% but increases time cost by 3.7 over the baseline method CoOp. In contrast, our Skip Tuning achieves both effectiveness and efficiency without compromise, further underscoring the superiority of our approach.